Solve the following pair of linear equations by the substitution method
 
s – t = 3
(s/3) + (t/2) = 6

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given,
s – t = 3
and
(s/3) + (t/2) = 6
are the two equations.
From 1st equation,
we get,
s = 3 + t ________________(1)
Now, substitute the value of s in second equation to get,
(3+t)/3 + (t/2) = 6
⇒ (2(3+t) + 3t )/6 = 6
⇒ (6+2t+3t)/6 = 6
⇒ (6+5t) = 36
⇒5t = 30

⇒t = 6
Now, substitute the value of t in equation (1)
s = 3 + 6 = 9
Therefore,
s = 9 and t = 6.

Thanks